P = (N/r)( 1 - (1 + r)-n )
P = (20000/0.05)( 1 - (1 + 0.05)^-20 )
= 400000(1 - .37689) = $249,244
This is the formula and the solution. Can someone PLEASE tell me how they did (1-(1+0.05)^-20) and reached the .37689.
I am desparate and no one has answered the last 2 questions I have posted. Am I posting in the wrong area?
: P = (N/r)( 1 - (1 + r)-n )
P = (20000/0.05)( 1 - (1 + 0.05)^-20 )
= 400000(1 - .37689) = $249,244
This is the formula and the solution. Can someone PLEASE tell me how they did (1-(1+0.05)^-20) and reached the .37689.
I am desparate and no one has answered the last 2 questions I have posted. Am I posting in the wrong area?

Answer by Earlsdon(3717)

Evaluate:
(1-(1+0.05)^(-20)) = (1-(1.05)^(-20))=(1-0.37689) = 0.62311 and...
(400000)*(0.62311) = 249244